OpenStudy (anonymous):

(7.1-(-1.4)-(0.2x4.5)-9.1/((-0.3)/0.2)

OpenStudy (kira_yamato):

(7.1-(-1.4)-(0.2x4.5)-9.1/((-0.3)/0.2) = 8.5-0.9-9.1/-1.5 = 13 2/3